OpenStudy (anonymous):

solve by substitution. 4x=6y+24. and 2x-3y=12

OpenStudy (alienium):

6y+24 = 4x 2x-3y=12 _______________ -4x+6y=-24 2x-3y=12 in second equation we localize x: 2x-3y=12 <==> x=(12+3y)/2 we substitute this wherever we have x in the first equation: -4[(12+3y)/2]+6y=-24 just solve this to get y and replace wherever you have y in the equation where we localized x and you get y

OpenStudy (anonymous):

how do you solve that though?

OpenStudy (anonymous):